OpenStudy (anonymous):

how do i simplify the square root of 1/11

OpenStudy (anonymous):

You don't?

OpenStudy (anonymous):

You can't because 11 is not a perfect square.

OpenStudy (anonymous):

if you want to write it in "simplest radical form" you write as follows \[\sqrt{\frac{1}{11}}=\frac{1}{\sqrt{11}}\times \frac{\sqrt{11}}{\sqrt{11}}\] \[=\frac{\sqrt{11}}{11}\]

OpenStudy (anonymous):

Or maybe 11^-1/2

OpenStudy (anonymous):

"simplest radical form" includes the the rule "no radicals in the denominator"

OpenStudy (anonymous):

Thank you satellite73

OpenStudy (anonymous):

so if you see \[\frac{1}{\sqrt{x}}\] you replace it by \[\frac{\sqrt{x}}{x}\]

OpenStudy (anonymous):

yw http://www.youtube.com/watch?v=K0DVk1S0bjk&feature=related rough rider!